OpenStudy (anonymous):

Which of the following is true? A. twenty-one is a composite number because one and twenty-one are its only factors. B. Thirteen is a prime number because it is only divisible by one and itself. C. sixteen is a prime number because it is divisible by 4. D. twenty-eight is a composite number because eighty-two is a prime number.

OpenStudy (anonymous):

composite numbers have factor or numbers that you can multiply to get to that number but prime numbers does not it only has 1 and itself

OpenStudy (anonymous):

The correct answer is B

OpenStudy (anonymous):

B. Thirteen is a prime number because it is only divisible by one and itself.

OpenStudy (anonymous):

@JAMG can you dumb it down a little ?

OpenStudy (anonymous):

Why is B correct ?

OpenStudy (anonymous):

composite = made of parts (factors) lets look at 28 1*28 = 28 2*14 = 28 4*7 = 28 2, 3, 5, 7, 11, 13 are all prime numbers because you can only get that number if you multiply that number times 1 there are not other numbers (factors)

OpenStudy (anonymous):

so B ? @JAMG <3

OpenStudy (anonymous):

yes, sorry I didn't realize I didn't answer the question

OpenStudy (anonymous):

lol its ok thank you :)